# LLM-Enhanced Multimodal Fusion for Cross-Domain Sequential Recommendation

### v2

@@ 1 Introduction
Sequential Recommendation (SR) has become a widely adopted approach for modeling dynamic user preferences by analyzing historical interaction sequences [1, 2, 3, 4]. While SR effectively predicts the next item within a single domain, it suffers from data sparsity and domain bias, where limited interaction data and overfitting to domain-specific patterns hinder generalization. Cross-Domain Sequential Recommendation (CDSR) [5, 6, 7, 8, 9, 10, 11] addresses these challenges by leveraging user behavior across multiple domains, enabling knowledge transfer and a more holistic modeling of user interests. However, existing CDSR methods still face three key limitations: (i) they primarily focus on intra-domain dependencies, underexploring inter-domain relationships; (ii) they insufficiently utilize multimodal information such as images and text; and (iii) even LLM-based approaches [12] rarely consider domain imbalance in cross-domain preference modeling. The difference between traditional CDSR and our approach is shown in Fig. 1.
@@
To address these gaps, we propose LLM-Enhanced Multimodal Fusion for Cross-Domain Sequential Recommendation (LLM-EMF), a novel framework that integrates prompt-designed Large Language Model (LLM) augmentation, multimodal fusion, and domain-balanced hierarchical attention. Unlike LLMRec [12], which augments textual features without explicitly mitigating domain imbalance, LLM-EMF generates domain-agnostic semantic attributes via LLM prompts to enhance cross-domain alignment. Compared with IFCDSR [13], which fuses image and ID embeddings but lacks explicit textual enrichment, our framework unifies visual, textual, and ID-based embeddings. In contrast to MAN [14], which applies mixed attention without balancing domain contributions, our hierarchical attention mechanism explicitly regulates domain influence, preventing high-frequency domains from dominating recommendations. To the best of our knowledge, LLM-EMF is the first CDSR framework that systematically integrates LLM-generated textual knowledge with multimodal fusion and explicit domain balance.
@@
    We conduct extensive experiments on four reorganized e-commerce datasets, demonstrating that LLM-EMF consistently outperforms recent state-of-the-art baselines [13, 14] across all evaluation metrics.
@@ 2.1 Sequential Recommendation
Sequential recommendation [15, 16, 17, 18] models user behavior as a sequence of time-sensitive items, aiming to predict the next item based on historical interactions. Early approaches, such as FPMC [19], used first-order Markov chains to model temporal dependencies. With growing behavioral complexity, deep learning models—RNNs [20, 21], CNNs [22], and attention mechanisms [23, 24, 25, 26, 27, 28]—were integrated into recommender systems [29, 30, 31, 32, 33], capturing long-term dependencies and dynamic preferences. Methods like SURGE [34] further reduced sequence complexity via metric learning for efficient large-scale retrieval. More recently, works such as DFAR [35] and DCN [36] have sought to capture deeper and more intricate relationships within sequential recommendation data using dual attention mechanisms and complex interaction modeling.
@@ 2.2 Cross-Domain Sequential Recommendation
Unlike single-domain sequential recommendation, which models behavior within one domain, cross-domain recommender systems [37, 38] address sparse data and cold-start issues by leveraging user behavior from multiple domains. Early methods [39] assumed auxiliary behaviors from other domains could enhance the target domain. Transfer learning [40] approaches (e.g., MiNet [41], CoNet [42], itemCST [43]) transfer learned representations to boost target-domain recommendations.
Industrial settings often require improving performance in all domains simultaneously, motivating dual learning [44, 45, 46, 47, 48, 49, 50, 51, 52], as in DTCDR and DDTCDR, to enable mutual enhancement via domain interaction modeling. However, many approaches overlook visual perceptual features and lack multimodal fusion with attention mechanisms, which can better focus on relevant signals from images, text, and other media, enriching user preference modeling across domains.
@@ 2.3 LLM-based Recommendation System
Large Language Models (LLMs) have gained significant traction in recommender systems due to their ability to leverage textual features of items and improve recommendation performance [53, 54, 55, 56, 57, 58]. Most existing LLM-based recommenders operate without fine-tuning, relying on pretrained knowledge to predict the next item [59, 60]. For example, CHAT-REC [61] uses ChatGPT to understand user preferences and deliver interactive, explainable recommendations. Similarly, GPT4Rec [62] employs GPT-2 to generate hypothetical ”search queries” from user history, which are then processed by the BM25 search engine to retrieve relevant items. Another line of research focuses on fine-tuning LLMs for specific recommendation tasks. TALLRec [63], for instance, applies instruction-tuning to decide whether an item should be recommended, while BIGRec [64] combines LLM-generated recommendations with collaborative filtering through an ensemble approach. Additionally, RecInterpreter [65] and LLaRA [66] treat ”sequential user behaviors” as a new modality for LLMs, aligning it with language representations. CoLLM [67] captures collaborative signals using an external model and integrates them into the LLM’s token embedding space, enabling the use of collaborative embeddings.
While some studies like TALLRec [63], BIGRec [64], and LLM-Rec [68] acknowledge the cross-domain potential of LLMs, there remains a gap in frameworks explicitly designed for CDSR that combine LLM-based textual enrichment with multimodal fusion and domain-balanced attention.
Compared with recent representative methods, our work addresses several gaps. First, unlike LLMRec [12], which augments item text but does not explicitly mitigate domain imbalance, our framework applies a prompt-designed LLM augmentation strategy to generate domain-agnostic attributes that improve cross-domain semantic alignment. Second, in contrast to IFCDSR [13], which fuses image and ID embeddings without enriched textual features, we unify visual, textual, and ID representations within a single multimodal fusion framework. Finally, whereas MAN [14] employs mixed attention without controlling domain contributions, our hierarchical attention mechanism explicitly balances the influence of each domain, preventing high-frequency domains from dominating recommendations.
@@ 3.3 Prompt with LLM
In the CDSR setting, we denote user–item interaction sequences as $X_{item}$ and their corresponding item metadata (e.g., titles and domains) as $T_{info}$. As illustrated in Fig. 3, for each item in the dataset, a pre-defined template is used as a language command for Deepseek-r1 [69] to generate additional contextual knowledge. This approach ensures the relevance, diversity, and richness of item information, which is critical for enhancing item representations across multiple domains.
@@
where $P_{n}$ represents the generated text prompt for enriching item representations, which is subsequently used as input for enhancing item embeddings. For each item, we design a unified template as the input command for Deepseek-r1 [69], leveraging the item’s title and description to generate additional contextual knowledge that further enriches the representation. This process requires only item-level metadata, making it both scalable and efficient.
@@ 3.5 Multiple Attention Mechanisms
In traditional sequential recommendation methods [29, 70], behavioral sequences from domains $X$ and $Y$ are often processed indiscriminately, leading to potential dominance by the domain with higher behavioral frequency. This imbalance can skew the representation of user preferences. To address this issue, we propose a hierarchical attentional framework that separately processes sequences $S^{X}$, $S^{Y}$, and $S^{X+Y}$, ensuring balanced representation across domains. Our approach integrates item identity (ID), visual features, and textual embeddings to model user preferences more comprehensively. As illustrated in Fig. 4, to model the interactions within each sequence more effectively, we incorporate a self-attention mechanism with query, key, and value (QKV) pairs. The attention mechanism allows the model to focus on relevant parts of the sequence when making predictions, capturing both local and global dependencies within the sequence.
@@
NCF-MLP [71] 4.49 3.94 4.51 2.18 1.57 2.03
$\pi$-Net [5] 7.68 7.32 8.13 3.53 2.98 3.73
GRU4Rec [29] 5.79 5.48 6.13 3.06 2.55 3.10
SASRec [70] 7.30 6.90 7.79 3.79 3.35 3.93
SR-GNN [72] 7.84 7.58 8.35 4.01 3.47 4.13
PSJNet [73] 8.33 8.07 8.77 4.10 3.68 4.32
MIFN [74] 8.55 8.28 9.01 4.09 3.57 4.29
Tri-CDR [75] 8.35 8.18 8.89 4.29 3.63 4.33
MAN [14] 8.65 8.42 9.41 4.78 4.47 5.05
LLMRec [12] 9.05 8.86 10.03 5.02 4.73 5.34
IFCDSR [13] 9.05 8.85 9.92 4.95 4.56 5.24
@@
NCF-MLP [71] 3.05 2.26 2.96 1.43 1.06 1.26
$\pi$-Net [5] 4.16 3.72 4.17 2.17 1.84 2.03
GRU4Rec [29] 3.83 3.14 3.73 1.68 1.34 1.52
SASRec [70] 3.79 3.23 3.69 1.81 1.41 1.71
SR-GNN [72] 3.85 3.27 3.78 1.78 1.40 1.66
PSJNet [73] 4.63 4.06 4.76 2.44 2.07 2.35
MIFN [74] 5.05 4.21 5.20 2.51 2.12 2.31
MAN [14] 5.95 5.06 5.45 2.64 2.34 2.57
Tri-CDR [75] 5.15 4.62 5.05 2.32 2.08 2.22
LLMRec [12] 6.21 5.29 5.71 2.80 2.48 2.74
IFCDSR [13] 6.08 5.02 5.86 2.75 2.37 2.65
@@ 4 Experiments
In this section, we introduce the experimental setup, providing detailed descriptions of the dataset, evaluation metrics, and implementation specifics. We then compare our method with state-of-the-art approaches on the Amazon dataset [76]. Finally, we conduct ablation studies to validate the effectiveness of the proposed method.
@@
Dataset and Evaluated Metric. Our experiments are conducted on the Amazon dataset [76], which serves as the foundation for constructing cross-domain sequential recommendation (CDSR) scenarios. The dataset offers a rich collection of user-item interactions across a variety of domains, making it ideal for modeling cross-domain behaviors. Building on previous works [5, 74], we select four distinct domains to create two CDSR scenarios for our experiments: “Food-Kitchen” and “Movie-Book.” These domains were chosen to reflect real-world applications where users frequently engage with multiple categories, allowing us to investigate the efficacy of cross-domain knowledge transfer in sequential recommendation settings. The process of preparing the data begins by extracting users who have interacted with items in both domains. To ensure meaningful analysis, we filter out users and items with fewer than 10 interactions, focusing only on those with sufficient behavioral data. This filtering step is crucial for ensuring the robustness of our results, as sparse data can lead to unreliable recommendations. Furthermore, in order to meet the sequential constraints necessary for our study, we retain cross-domain interaction sequences that contain at least three items from each domain. This ensures that each sequence provides enough context for learning temporal dependencies and domain transitions. For the training/validation/test split, we adopt a temporal strategy to preserve the sequential nature of the data. Specifically, we divide the latest interaction sequences into validation and test sets, while the remaining sequences are allocated to the training set. This approach helps simulate a real-world scenario, where past user behavior informs future recommendations. The statistics of our refined datasets for the CDSR scenarios, including the number of users, items, and interactions in each domain, are summarized in Tab. 1.
@@
To evaluate the performance of our model, we adopt two widely used evaluation metrics in recommendation systems: the Mean Reciprocal Rank (MRR) metric [77], which measures the rank of the first relevant item, and the Normalized Discounted Cumulative Gain (NDCG5, 10) metric  [78]. MRR is a metric that assesses the ranking quality by calculating the average of the reciprocal ranks of the first relevant item in each user’s recommendation list. NDCG, on the other hand, takes into account both the position and relevance of the recommended items, making it particularly suitable for evaluating ranked lists. These metrics are standard in the evaluation of recommendation systems, providing a comprehensive measure of recommendation quality. The results obtained from these metrics allow us to compare our model’s performance across different CDSR scenarios and demonstrate its effectiveness in capturing cross-domain interactions.
@@
Implementation Details. For fair comparisons, we adopt the same hyperparameter settings as in previous works [74]. We conduct experiments with various hyperparameter settings, as shown in Figure 5. Specifically, we set the learnable embedding size $q$ to 256, the CLIP-based image and text embedding size $e$ to 512, and the mini-batch size to 256, with a dropout rate of 0.3. The values of $\lambda_{1}$ and $\lambda_{2}$ are set to 0.3 and 0.1, respectively. The $L_{2}$ regularization coefficient is selected from $\{0.00001,0.00005,0.0001,0.0005,0.001\}$, and the learning rate is chosen from $\{0.0001,0.0005,0.001,0.005,0.01\}$. Training is conducted for 100 epochs using an NVIDIA 4090 GPU, with the Adam optimizer [79] to update all parameters. To ensure robustness and reproducibility, we perform a grid search over the hyperparameter space, evaluating the model’s performance on a validation set at each step. The selected hyperparameters are those that yield the best validation performance. Additionally, we employ early stopping with a patience of 10 epochs to prevent overfitting, monitoring the validation loss during training. This approach ensures that the model generalizes well to unseen data while maintaining computational efficiency. The use of the Adam optimizer [79] is particularly advantageous due to its adaptive learning rate mechanism, which helps stabilize training and accelerate convergence. Furthermore, the NVIDIA 4090 GPU provides the necessary computational power to handle the large-scale datasets and complex model architectures efficiently. These settings collectively ensure that our experiments are conducted under optimal conditions, enabling meaningful comparisons with state-of-the-art methods in the field.

### v3

# arXiv:2506.17966v3[cs.IR] 27 Feb 2026

@@ 1 Introduction
Sequential Recommendation (SR) has become a widely adopted approach for modeling dynamic user preferences by analyzing historical interaction sequences Shani et al. [2002], He et al. [2017a], He and McAuley [2016], Li et al. [2017]. While SR effectively predicts the next item within a single domain, it suffers from data sparsity and domain bias, where limited interaction data and overfitting to domain-specific patterns hinder generalization. Cross-Domain Sequential Recommendation (CDSR) Ma et al. [2019], Zhuang et al. [2020], Guo et al. [2021], Lei et al. [2021], Cao et al. [2022], Li et al. [2022], Jiang et al. [2023] addresses these challenges by leveraging user behavior across multiple domains, enabling knowledge transfer and a more holistic modeling of user interests. However, existing CDSR methods still face three key limitations: (i) they primarily focus on intra-domain dependencies, underexploring inter-domain relationships; (ii) they insufficiently utilize multimodal information such as images and text; and (iii) even LLM-based approaches Wei et al. [2024] rarely consider domain imbalance in cross-domain preference modeling. The difference between traditional CDSR and our approach is shown in Fig. 1.
@@
To address these gaps, we propose LLM-Enhanced Multimodal Fusion for Cross-Domain Sequential Recommendation (LLM-EMF), a novel framework that integrates prompt-designed Large Language Model (LLM) augmentation, multimodal fusion, and domain-balanced hierarchical attention. Unlike LLMRec Wei et al. [2024], which augments textual features without explicitly mitigating domain imbalance, LLM-EMF generates domain-agnostic semantic attributes via LLM prompts to enhance cross-domain alignment. Compared with IFCDSR Wu et al. [2025b], which fuses image and ID embeddings but lacks explicit textual enrichment, our framework unifies visual, textual, and ID-based embeddings. In contrast to MAN Lin et al. [2024], which applies mixed attention without balancing domain contributions, our hierarchical attention mechanism explicitly regulates domain influence, preventing high-frequency domains from dominating recommendations. To the best of our knowledge, LLM-EMF is the first CDSR framework that systematically integrates LLM-generated textual knowledge with multimodal fusion and explicit domain balance.
@@
    We conduct extensive experiments on four reorganized e-commerce datasets, demonstrating that LLM-EMF consistently outperforms recent state-of-the-art baselines Wu et al. [2025b], Lin et al. [2024] across all evaluation metrics.
@@ 2.1 Sequential Recommendation
Sequential recommendation Cai et al. [2024], Wang et al. [2019], Liang et al. [2025], Wu et al. [2025a] models user behavior as a sequence of time-sensitive items, aiming to predict the next item based on historical interactions. Early approaches, such as FPMC Rendle et al. [2010], used first-order Markov chains to model temporal dependencies. With growing behavioral complexity, deep learning models—RNNs Chung et al. [2014], Hochreiter and Schmidhuber [1997], CNNs LeCun et al. [1995], and attention mechanisms Vaswani et al. [2017]—were integrated into recommender systems Hidasi et al. [2015], Zhou et al. [2019], Tang and Wang [2018], Kang and McAuley [2018a], Zhou et al. [2018], capturing long-term dependencies and dynamic preferences. Methods like SURGE Chang et al. [2021] further reduced sequence complexity via metric learning for efficient large-scale retrieval. More recently, works such as DFAR Lin et al. [2023] and DCN Lin et al. [2022] have sought to capture deeper and more intricate relationships within sequential recommendation data using dual attention mechanisms and complex interaction modeling.
@@ 2.2 Cross-Domain Sequential Recommendation
Unlike single-domain sequential recommendation, which models behavior within one domain, cross-domain recommender systems Xiao et al. [2025], Cantador et al. [2015] address sparse data and cold-start issues by leveraging user behavior from multiple domains. Early methods Singh and Gordon [2008] assumed auxiliary behaviors from other domains could enhance the target domain. Transfer learning Pan and Yang [2009] approaches (e.g., MiNet Ouyang et al. [2020], CoNet Hu et al. [2018], itemCST Pan et al. [2010]) transfer learned representations to boost target-domain recommendations.
Industrial settings often require improving performance in all domains simultaneously, motivating dual learning Long et al. [2012], He et al. [2016], Hou et al. [2024], Li et al. [2025, 2024b, 2024a, 2024c], Liu et al. [2024, 2025], as in DTCDR and DDTCDR, to enable mutual enhancement via domain interaction modeling. However, many approaches overlook visual perceptual features and lack multimodal fusion with attention mechanisms, which can better focus on relevant signals from images, text, and other media, enriching user preference modeling across domains.
@@ 2.3 LLM-based Recommendation System
Large Language Models (LLMs) have gained significant traction in recommender systems due to their ability to leverage textual features of items and improve recommendation performance Geng et al. [2022], Liu et al. [2023], Wu et al. [2023], Qiu et al. [2024, 2025b, 2025a]. Most existing LLM-based recommenders operate without fine-tuning, relying on pretrained knowledge to predict the next item Sun et al. [2023], Wang et al. [2023]. For example, CHAT-REC Gao et al. [2023] uses ChatGPT to understand user preferences and deliver interactive, explainable recommendations. Similarly, GPT4Rec Li et al. [2023] employs GPT-2 to generate hypothetical ”search queries” from user history, which are then processed by the BM25 search engine to retrieve relevant items. Another line of research focuses on fine-tuning LLMs for specific recommendation tasks. TALLRec Bao et al. [2023b], for instance, applies instruction-tuning to decide whether an item should be recommended, while BIGRec Bao et al. [2023a] combines LLM-generated recommendations with collaborative filtering through an ensemble approach. Additionally, RecInterpreter Yang et al. [2023] and LLaRA Liao et al. [2023] treat ”sequential user behaviors” as a new modality for LLMs, aligning it with language representations. CoLLM Zhang et al. [2025] captures collaborative signals using an external model and integrates them into the LLM’s token embedding space, enabling the use of collaborative embeddings.
While some studies like TALLRec Bao et al. [2023b], BIGRec Bao et al. [2023a], and LLM-Rec Tang et al. [2023] acknowledge the cross-domain potential of LLMs, there remains a gap in frameworks explicitly designed for CDSR that combine LLM-based textual enrichment with multimodal fusion and domain-balanced attention.
Compared with recent representative methods, our work addresses several gaps. First, unlike LLMRec Wei et al. [2024], which augments item text but does not explicitly mitigate domain imbalance, our framework applies a prompt-designed LLM augmentation strategy to generate domain-agnostic attributes that improve cross-domain semantic alignment. Second, in contrast to IFCDSR Wu et al. [2025b], which fuses image and ID embeddings without enriched textual features, we unify visual, textual, and ID representations within a single multimodal fusion framework. Finally, whereas MAN Lin et al. [2024] employs mixed attention without controlling domain contributions, our hierarchical attention mechanism explicitly balances the influence of each domain, preventing high-frequency domains from dominating recommendations.
@@ 3.3 Prompt with LLM
In the CDSR setting, we denote user–item interaction sequences as $X_{item}$ and their corresponding item metadata (e.g., titles and domains) as $T_{info}$. As illustrated in Fig. 3, for each item in the dataset, a pre-defined template is used as a language command for Deepseek-r1 Guo et al. [2025] to generate additional contextual knowledge. This approach ensures the relevance, diversity, and richness of item information, which is critical for enhancing item representations across multiple domains.
@@
where $P_{n}$ represents the generated text prompt for enriching item representations, which is subsequently used as input for enhancing item embeddings. For each item, we design a unified template as the input command for Deepseek-r1 Guo et al. [2025], leveraging the item’s title and description to generate additional contextual knowledge that further enriches the representation. This process requires only item-level metadata, making it both scalable and efficient.
@@ 3.5 Multiple Attention Mechanisms
In traditional sequential recommendation methods Hidasi et al. [2015], Kang and McAuley [2018b], behavioral sequences from domains $X$ and $Y$ are often processed indiscriminately, leading to potential dominance by the domain with higher behavioral frequency. This imbalance can skew the representation of user preferences. To address this issue, we propose a hierarchical attentional framework that separately processes sequences $S^{X}$, $S^{Y}$, and $S^{X+Y}$, ensuring balanced representation across domains. Our approach integrates item identity (ID), visual features, and textual embeddings to model user preferences more comprehensively. As illustrated in Fig. 4, to model the interactions within each sequence more effectively, we incorporate a self-attention mechanism with query, key, and value (QKV) pairs. The attention mechanism allows the model to focus on relevant parts of the sequence when making predictions, capturing both local and global dependencies within the sequence.
@@
NCF-MLP He et al. [2017b] 4.49 3.94 4.51 2.18 1.57 2.03
$\pi$-Net Ma et al. [2019] 7.68 7.32 8.13 3.53 2.98 3.73
GRU4Rec Hidasi et al. [2015] 5.79 5.48 6.13 3.06 2.55 3.10
SASRec Kang and McAuley [2018b] 7.30 6.90 7.79 3.79 3.35 3.93
SR-GNN Wu et al. [2019] 7.84 7.58 8.35 4.01 3.47 4.13
PSJNet Sun et al. [2022] 8.33 8.07 8.77 4.10 3.68 4.32
MIFN Ma et al. [2022] 8.55 8.28 9.01 4.09 3.57 4.29
Tri-CDR Ma et al. [2024] 8.35 8.18 8.89 4.29 3.63 4.33
MAN Lin et al. [2024] 8.65 8.42 9.41 4.78 4.47 5.05
LLMRec Wei et al. [2024] 9.05 8.86 10.03 5.02 4.73 5.34
IFCDSR Wu et al. [2025b] 9.05 8.85 9.92 4.95 4.56 5.24
@@
NCF-MLP He et al. [2017b] 3.05 2.26 2.96 1.43 1.06 1.26
$\pi$-Net Ma et al. [2019] 4.16 3.72 4.17 2.17 1.84 2.03
GRU4Rec Hidasi et al. [2015] 3.83 3.14 3.73 1.68 1.34 1.52
SASRec Kang and McAuley [2018b] 3.79 3.23 3.69 1.81 1.41 1.71
SR-GNN Wu et al. [2019] 3.85 3.27 3.78 1.78 1.40 1.66
PSJNet Sun et al. [2022] 4.63 4.06 4.76 2.44 2.07 2.35
MIFN Ma et al. [2022] 5.05 4.21 5.20 2.51 2.12 2.31
MAN Lin et al. [2024] 5.95 5.06 5.45 2.64 2.34 2.57
Tri-CDR Ma et al. [2024] 5.15 4.62 5.05 2.32 2.08 2.22
LLMRec Wei et al. [2024] 6.21 5.29 5.71 2.80 2.48 2.74
IFCDSR Wu et al. [2025b] 6.08 5.02 5.86 2.75 2.37 2.65
@@ 4 Experiments
In this section, we introduce the experimental setup, providing detailed descriptions of the dataset, evaluation metrics, and implementation specifics. We then compare our method with state-of-the-art approaches on the Amazon dataset Wei et al. [2021]. Finally, we conduct ablation studies to validate the effectiveness of the proposed method.
@@
Dataset and Evaluated Metric. Our experiments are conducted on the Amazon dataset Wei et al. [2021], which serves as the foundation for constructing cross-domain sequential recommendation (CDSR) scenarios. The dataset offers a rich collection of user-item interactions across a variety of domains, making it ideal for modeling cross-domain behaviors. Building on previous works Ma et al. [2019, 2022], we select four distinct domains to create two CDSR scenarios for our experiments: “Food-Kitchen” and “Movie-Book.” These domains were chosen to reflect real-world applications where users frequently engage with multiple categories, allowing us to investigate the efficacy of cross-domain knowledge transfer in sequential recommendation settings. The process of preparing the data begins by extracting users who have interacted with items in both domains. To ensure meaningful analysis, we filter out users and items with fewer than 10 interactions, focusing only on those with sufficient behavioral data. This filtering step is crucial for ensuring the robustness of our results, as sparse data can lead to unreliable recommendations. Furthermore, in order to meet the sequential constraints necessary for our study, we retain cross-domain interaction sequences that contain at least three items from each domain. This ensures that each sequence provides enough context for learning temporal dependencies and domain transitions. For the training/validation/test split, we adopt a temporal strategy to preserve the sequential nature of the data. Specifically, we divide the latest interaction sequences into validation and test sets, while the remaining sequences are allocated to the training set. This approach helps simulate a real-world scenario, where past user behavior informs future recommendations. The statistics of our refined datasets for the CDSR scenarios, including the number of users, items, and interactions in each domain, are summarized in Tab. 1.
@@
To evaluate the performance of our model, we adopt two widely used evaluation metrics in recommendation systems: the Mean Reciprocal Rank (MRR) metric Voorhees [1999], which measures the rank of the first relevant item, and the Normalized Discounted Cumulative Gain (NDCG5, 10) metric  Järvelin and Kekäläinen [2002]. MRR is a metric that assesses the ranking quality by calculating the average of the reciprocal ranks of the first relevant item in each user’s recommendation list. NDCG, on the other hand, takes into account both the position and relevance of the recommended items, making it particularly suitable for evaluating ranked lists. These metrics are standard in the evaluation of recommendation systems, providing a comprehensive measure of recommendation quality. The results obtained from these metrics allow us to compare our model’s performance across different CDSR scenarios and demonstrate its effectiveness in capturing cross-domain interactions.
@@
Implementation Details. For fair comparisons, we adopt the same hyperparameter settings as in previous works Ma et al. [2022]. We conduct experiments with various hyperparameter settings, as shown in Figure 5. Specifically, we set the learnable embedding size $q$ to 256, the CLIP-based image and text embedding size $e$ to 512, and the mini-batch size to 256, with a dropout rate of 0.3. The values of $\lambda_{1}$ and $\lambda_{2}$ are set to 0.3 and 0.1, respectively. The $L_{2}$ regularization coefficient is selected from $\{0.00001,0.00005,0.0001,0.0005,0.001\}$, and the learning rate is chosen from $\{0.0001,0.0005,0.001,0.005,0.01\}$. Training is conducted for 100 epochs using an NVIDIA 4090 GPU, with the Adam optimizer Kingma and Ba [2015] to update all parameters. To ensure robustness and reproducibility, we perform a grid search over the hyperparameter space, evaluating the model’s performance on a validation set at each step. The selected hyperparameters are those that yield the best validation performance. Additionally, we employ early stopping with a patience of 10 epochs to prevent overfitting, monitoring the validation loss during training. This approach ensures that the model generalizes well to unseen data while maintaining computational efficiency. The use of the Adam optimizer Kingma and Ba [2015] is particularly advantageous due to its adaptive learning rate mechanism, which helps stabilize training and accelerate convergence. Furthermore, the NVIDIA 4090 GPU provides the necessary computational power to handle the large-scale datasets and complex model architectures efficiently. These settings collectively ensure that our experiments are conducted under optimal conditions, enabling meaningful comparisons with state-of-the-art methods in the field.